DQuestion 1 1 pts You expect to receive $4197 in 19 years. How much is this amount worth to you today given your required interst rate is 5.8%?
Answer #1

The amount worth today for amount receivable in future is the present value of the amount receivable

Present Value factor

= 1 / (1 + r) ^ n

Where,

r = Rate of interest = 5.8% or 0.058

n = Years = 19

So, PV Factor

= 1 / 1.058^19

= 1 / 2.918957

= 0.342588

So, Present Value

= Future value receivable x Present value factor

= $4,197 x 0.342588

= $1,437.84